OpenStudy (anonymous):

2x-5y=-7 5x-3y=11 solve by elimination

OpenStudy (anonymous):

@Peaches15

OpenStudy (anonymous):

i did what u showed me and i stuck with 6x-2y=4 @Peaches15

OpenStudy (anonymous):

if you would want to solve for y it would be 2x-5y=-7}1 5x-3y=11}2 so you would have to multiply the first equation by 5 and the second equation by 2 because it is the lowest common multiple 10x-25y=-35 10x-6y=22 now subtract because you want to solve for y 19y=-57 y=3 now substitute y in to any problem to solve for x 2x-5(3)=-7 2x-15=-7 2x=8 x=4 (4,3) @Pandabear78963 I like these questions der easy if you have more of dezz ask me

OpenStudy (anonymous):

no prob dont worry i will cuz when i @ u......u come asap :) @Peaches15

OpenStudy (anonymous):

can u help me with my math its not dat hard

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill try my best @Peaches15